SILICON-PHOSPHORUS BONDING: LABORATORY DETECTION OF HPSiH_{2} EMPLOYING HIGH RESOLUTION MICROWAVE SPECTROSCOPY

\underline{VALERIO LATTANZI}, M.C. McCARTHY, P. THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, and School of Engineering and Applied Science, Harvard University, Cambridge, MA 02138; and SVEN THORWIRTH, Max-Planck-Institut für Radioastronomie, Bonn, Germany, and I. Physikalisches Institut, Universität zu Köln, Germany.

HPSiH_{2}, the ground state isomer on the H_{3}SiP potential energy surface, has been detected by means of Fabry-Pérot FT microwave spectroscopy. The laboratory search has been guided by theoretical structure calculations performed at the CCSD(T)/cc-pwCVQZ level of theory corrected for zero-point vibrational effects at the CCSD(T)/cc-pV(T+d)Z level. A mixture of silane and phosphine in a discharge supersonic molecular beam has been used to produce the new species, allowing the detection of the three lowest K_{a}=0 rotational transitions. The discovery has been confirmed by successful identification of the same transitions of HP^{29}SiH_{2}, HP^{30}SiH_{2}, and DPSiD_{2}, at precisely the expected frequency shifts. The presence of other Si and some P bearing molecules in astronomical sources suggests, that this molecule is a plausible candidate for radio astronomical detection.